40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 16:29:40 2020, gautam, Update, Electronics, c1psl VME crate removed, Acro-crate installed c1psl.pdf
    Reply  Tue Mar 3 17:20:14 2020, gautam, Update, Electronics, More cabling removed 
       Reply  Wed Mar 4 16:18:31 2020, gautam, Update, Electronics, More cabling removed 1X2EuroBefore.JPGIOO.png
          Reply  Wed Mar 4 21:02:49 2020, Koji, Update, Electronics, More cabling removed 
    Reply  Tue Mar 3 17:59:33 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
       Reply  Thu Mar 5 15:03:48 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
          Reply  Thu Mar 5 19:45:23 2020, Jon, Summary, PSL, C1PSL in-situ test results c1psl_feedthrough_wiring_-_By_Connector_(3).pdf
             Reply  Wed Mar 11 18:12:53 2020, gautam, Summary, PSL, WFS Demod board modifications D980233-B_Mar2020Mods.pdf
Message ID: 15256     Entry time: Thu Mar 5 19:45:23 2020     In reply to: 15255     Reply to this: 15266
Author: Jon 
Type: Summary 
Category: PSL 
Subject: C1PSL in-situ test results 

We've completed almost all of the in-situ testing of the c1psl channels. During this process, we identified several channels which needed to be rewired to different Acromags (BIO sinking v. sourcing). We also elected to change the connector type of a few channels for practical advantages. Those modifications and other issues found during testing are detailed below. Also attached are the updated channel assignments, with a column indicating the in-situ testing status of each channel.

Post-installation modifications

  • All four channels connected to the sourcing BIO module were found to in fact require sinking I/O. They were reassigned to sinking BIO modules. Affected channels:
    • C1:PSL-FSS_FASTSWEEP
    • C1:PSL-FSS_SW1
    • C1:PSL-FSS_SW2
    • C1:PSL-PSL_Shutter
  • Added a new AI channel:
    • C1:PSL-FSS_MIXERM
  • Removed an unneccessary AI channel:
    • C1:PSL-FSS_LODET
  • Moved two AI channels from BNC connectors to a new Dsub connector (labelled DB25M-2 in the spreadsheet).
    • C1:PSL-FSS_RCTEMP
    • C1:PSL-FSS_RMTEMP_VOLTS

Issues identified during testing

  • Digital calibration. The following channels work, but we need to verify their EPICS calibration parameters (EGUF/EGUL):
    • C1:PSL-FSS_FASTGAIN
    • C1:PSL-FSS_FAST
    • C1:PSL-PMC_RFADJ
    • C1:PSL-PMC_MODET
  • IMC servo board. The Acromag channels themselves were found to work, but the linearity of the mbbo gain stages are in question (i.e., a potential problem with the board). GV is currently testing the servo board.
  • PSL QPD board apears to be dead. We connected a scope directly to the test points on the board and measured a high level of noise and no signal (for all four of the QPD channels). I understand this QPD has not been used in some time, so it may not have been noticed before.
  • WFS DC channels are saturating when the IMC is unlocked. The acceptance range of the Acromag ADC is only +/-10 V, but we measured sensor voltages as high as ~14 V. It appears that the old ADCs were somehow accepting a range of 0 to +20 V instead of -10 to +10 V. However, the Acromags do not support the input range 0-20 V. Since SNR is not critical for these channels (they're used only for initial alignment), I propose we simply install a voltage divider inside the chassis, just before the Acromag, for each of these signals.
Attachment 1: c1psl_feedthrough_wiring_-_By_Connector_(3).pdf  119 kB  | Show | Show all
ELOG V3.1.3-